OpenStudy (anonymous):

Brett has been studying a type of bacteria that doubles every month. Originally, there were 5 bacterial cells. He wants to know how many there will be after 42 months. Which equation should he use? A. a42 = 5(2)42 − 1 B. a42 = 5(2)(42) C. a42 = 5(2)42 D. a42 = 5(2)(42 − 1)

zepdrix (zepdrix):

At the very start, when no months have passed, we have 5 bacteria cells. Let's call this \(\Large\rm a_0\), the amount \(\Large\rm a\). The subscript 0 tells us how many months have passed. After one month, \(\Large\rm a_1=5(2)\) we double the 5. After two months, \(\Large\rm a_2=5(2)^2\) we double the 5, then we double again. After n months have passed, \(\Large\rm a_n=5(2)^n\) we double n consecutive number of times. So what about after 42 months, what form do you think it will have? :) Pssst! Plug in 42 for n.